# SEIR models with host heterogeneity: theoretical aspects and applications to seasonal influenza dynamics

### v1

@@ Abstract
Population heterogeneity is a key factor in epidemic dynamics, influencing both transmission and final epidemic size. While heterogeneity is often modelled through age structure, spatial location, or contact patterns, differences in host susceptibility have recently gained attention, particularly during the COVID-19 pandemic. Building on the framework of Diekmann and Inaba (Journal of Mathematical Biology, 2023), we focus on the special case of SEIR models, which are widely used for influenza and other respiratory infections. We derive the model equations under two distinct assumptions linking susceptibility and infectiousness. Analytical results show that heterogeneity in susceptibility reduces the epidemic final size compared to homogeneous models with the same basic reproduction number $\mathcal{R}_{0}$. In the case of gamma-distributed susceptibility, we obtain stronger results on the epidemic final size. The resulting model captures population heterogeneity through a single parameter, which makes it practical for fitting epidemic data. We illustrate its use by applying it to seasonal influenza in Italy.
@@ 1 Introduction
Population heterogeneity has been recognised as a fundamental determinant of epidemic dynamics. Heterogeneities may be due to age [AndMay, Rost2020] or spatial location [Longini_old], and in this context, the structure of contact matrices [Polymod] or connections between nodes [Colizza2007] play an important role; heterogeneity in the number of contacts has been considered especially in the context of sexually-transmitted infections [Hym88], in which case the type of mixing (assortative or proportionate) significantly affects the epidemic dynamics. A different kind of heterogeneity, differences in susceptibility, i.e. in the probability of getting infected when contacted by an infectious individual, has been proposed by several authors [BrittonBallTrapman2020] during the Covid epidemic and systematically developed by [Novozhilov2008] and, more recently, by [DiekmannInaba2023] (see also [Bootsma2023] and [Inaba2025]). A somewhat similar model had been studied by [Hyman2005], who analysed the long-term dynamics of an endemic SIR model in which the population is divided into a discrete number of groups differing in susceptibility to infection. Different approaches to incorporate heterogeneity have been proposed by [Rose2021] and by [Berestycki2023].
[DiekmannInaba2023] developed the theory for a general type of epidemic in which infectiousness depends, in an arbitrary way, on the time since infection, and the model is set as a renewal equation; later, they specialise the resulting equation for the case in which the model can be described by a system of ODEs. Here, we consider only the case of SEIR models, which are widely used to describe the dynamics of influenza epidemics, as well as several other infectious diseases.
We begin by deriving the equations for this special case, following [DiekmannInaba2023], under two alternative assumptions on the relationship between susceptibility and infectiousness. By concentrating on this specific setting, the derivation becomes more transparent, even though the resulting systems have already been presented in [DiekmannInaba2023]. We then study the epidemic final size for the resulting models; we first show, similarly to [Novozhilov2008], that the presence of heterogeneity always decreases the epidemic final size relative to a homogeneous model with the same value of $\mathcal{R}_{0}$. We then consider the case in [DiekmannInaba2023] where susceptibility heterogeneity follows a gamma distribution and show that greater variance in susceptibility lowers the attack ratio. Lastly, we compare models with and without susceptibility heterogeneity by fitting them to Italian seasonal influenza data from 2014-15 to 2023-24
@@ 2 Model construction
The feature $x$ could depend on genetic factors, immunological history, or possibly behaviour. Still, it is assumed that it is fixed, at least as long as an epidemic outbreak is concerned, and that its distribution in the population has density $\phi(x)$ (a more general distribution, allowing also for a discrete part, is considered in [DiekmannInaba2023], but here we stick to the simplest case).
Following [DiekmannInaba2023], we can set $a(x)=x$ (one needs only to redefine the feature) and assume that its mean is 1, i.e., $\int_{0}^{\infty}x\phi(x)\,\textnormal{d}x=1$, by redefining $F(t)$ appropriately.
@@
The subsequent elaborations depend on the choice of the function $c(x)$; following [DiekmannInaba2023], we consider the cases $c(x)\equiv\beta$ (infectiousness independent of susceptibility) and $c(x)=\beta x$ (perfect correlation between susceptibility and infectiousness).
@@
While (8) is a closed system, it may be convenient to consider also the variables $E(t)$ and $V(t)$, in order to compare the model results with data. One has
@@
Although the function $\Phi_{k}$ cannot generally be explicitly computed (but see below the case of the gamma distribution for $x$), there is no problem in numerically integrating system (7).
@@ 3 Basic properties
In the case of correlated infectiousness and susceptibilities, $\Phi_{2}(1)$ involves the variance of the susceptibility $X$, since $\Phi_{2}(1)=1+{Var(x)}$. The resulting formula is similar to that obtained with the proportionate mixing in the case of varying contact rates [AndMay1986].
@@
The proof can be found in Appendix A. In biological terms, Theorem 3.1 shows that with any distribution of heterogeneity in susceptibility, and with both models, the susceptible fraction at the end of an epidemic is larger than with homogeneous susceptibility; in other words, the attack ratio (or epidemic final size) will be smaller. This is similar to the results obtained by [Novozhilov2008].
@@ 4.1 The case of gamma-distributed susceptibility
Again following [DiekmannInaba2023], we assume that $X$ follows the gamma distribution with mean 1 and variance $1/p$, i.e.,
@@ 4.2 Beta-distributed susceptibility
Here we assume that $X$ follows the Beta distribution scaled by a factor $L>1$, so that its support becomes $(0,L)$. Then
@@
This expression can be rewritten in terms of Kummer’s function $M(a,b,z)$ [Abramowitz, p. 505], also known as the hypergeometric confluent function $_{1}F_{1}(a,b,z)$ whose numerical computation is provided in several software packages (e.g. Wolfram-Alpha, Python, or Matlab), making computations much faster. Precisely, one has
@@ 4.3 Numerical illustrations
Figure 2: Plots showing proportion of $E(t)+I(t)$ (above) and $R(t)$ (below) for differently distributed susceptibility over time (all with $\mathcal{R}_{0}=2$). On the left panels, plots are shown with infectiousness independent of susceptibility; on the right panels, one can see plots with directly correlated infectiousness and susceptibility. The black curve represents the homogeneous SEIR model, coloured solid curves represent Gamma distributions with different variance (see legend), coloured dashed curves represent Beta-distributed heterogeneity, with the same variance $(Var(X)=1/3)$ as the light-blue, solid curve.
It can be seen that the homogeneous SEIR model yields the largest infection peak and the largest final size. Moreover, as the variance of the Gamma distribution increases, the final size systematically decreases, both when susceptibility and infectiousness are independent and when they are correlated. This dependence of the final size on will be proved in the next section. One can also observe that different Beta distributions with the same variance produce very similar final epidemic sizes, both among themselves and relative to the Gamma distribution with the same variance. In this sense, the epidemic dynamics appear to depend primarily on the variance of the susceptibility distribution rather than on its specific shape: although the three distributions shown in the right panel of Figure 1 differ markedly in form, they generate very similar trajectories and final epidemic sizes.
5 Final size equation for the Gamma distribution
@@ 5.1 The case of independent infectiousness and susceptibility
Since the variance of the Gamma distribution decreases with increasing $p$, this Theorem shows that the epidemic final size $(1-S_{\infty}(p)$ decreases with increasing variance. This result had been essentially proved by [Novozhilov2008] with completely different notation and methods.
@@ 5.2 The case of correlated susceptibility and infectiousness
is constant along the solutions of (15). Moreover, it is not difficult to show that necessarily
@@
Figure 3: Susceptible fraction at the end of an epidemic for model (7), plotted as a function of the variance. Curves are shown for the Gamma distribution and the Beta distributions with either the parameter choices in (19) or (20). In all cases, $\mathcal{R}_{0}=1.5$.
In Figure 3, we show the numerically computed dependence of the final susceptible fraction on the variance for the Gamma distribution and for the Beta distribution under the parametrisations (19) and (20). Overall, the final epidemic size obtained from the Beta distribution is close to that produced by a Gamma distribution with the same variance, though small differences persist. In particular, the Beta distributions tend to yield slightly larger values of $S_{\infty}$, and therefore a lower attack ratio. When comparing Beta and Gamma distributions, or Beta distributions with different parametrisations, it is possible to find pairs of distributions for which the one with the lower variance also has the lower attack ratio, in line with the observation of [Novozhilov2008] that mean and variance do not fully characterise the behaviour of heterogeneous models. Nevertheless, within each fixed Beta parametrisation, increasing the variance consistently leads to a lower attack ratio.
@@ 6.1 Data description
The data used in this study were obtained from RespiVirNet [ISS], which provides weekly, age-stratified influenza-like illness (ILI) incidence rates from the national sentinel surveillance system. Raw records include ILI aggregated case counts (and number of patients in the sentinel cohorts) from the 2014-15 season onwards, reported by age groups (0–4, 5–14, 15–64, 65 and above) and week (from Week 46 to Week 17 of the subsequent year). Furthermore, data include, for each week, the fraction of virological samples from the sentinel network that are positive to the different influenza strains; these are age-stratified only in recent years, thus, for consistency, we did not use this information.
@@
  • $ur_{a}$ is the estimate of age-dependent underreporting ratio [Trentini2022].
The influenza seasons considered range from 2014-15 to 2023-24. However, the seasons 2020-21, 2021-22 and 2022-23 have been excluded from the analysis since the lockdowns and other protective measures strongly affected the dynamics (no influenza cases were recorded in 2020-21, very few in 2021-22, while 2022-23 had an anomalous pattern, incompatible with a SEIR model with constant parameters).
@@ 6.2 Fitting methods
For the model (14) with Gamma-distributed susceptibility, we considered as unknown parameters $\beta$, $p$, and the initial condition $E(0)=I(0)$, where 0 is the first week of a given epidemic season; at $t=0$ everybody else was assumed to be susceptible, as nobody was considered completely immune to the circulating strain at the season start. In the case of the homogeneous SEIR model (formally, it can be seen as model (14) with $p=\infty$), we instead fitted $\beta$ and the initial conditions $E(0)=I(0)$ and $R(0)$. For both models, we set $\alpha=\gamma=4$ (weeks-1), corresponding to an average length of generation time of 3.5 days [Vink2014].

### v2

@@ Abstract
Population heterogeneity is a key factor in epidemic dynamics, influencing both transmission and final epidemic size. While heterogeneity is often modelled through age structure, spatial location, or contact patterns, differences in host susceptibility have recently gained attention, particularly during the COVID-19 pandemic. Building on the framework of Diekmann and Inaba (Journal of Mathematical Biology, 2023), we focus on the special case of SEIR epidemic models, assuming that at the epidemic start there is no pre-existing immunity. Under two distinct assumptions linking susceptibility and infectiousness, one obtains a closed system of 3 ODEs, which can be easily simulated and for which some analytical results are obtained. In particular, we proved that heterogeneity in susceptibility reduces the epidemic final size compared to homogeneous models with the same basic reproduction number $R_{0}$. We specialised in the case where susceptibility is distributed according to a gamma or extended Beta distribution, showing how the epidemic final size depends on the variance of the distribution. In the case of a gamma-distributed susceptibility, the resulting model consists of a system of ODEs with just one parameter more than the classical SEIR model; this makes it practical for fitting epidemic data. We illustrate its use by fitting data on seasonal influenza in Italy, and comparing the results to those obtained with simple SEIR models with pre-existing immunity.
@@ 1 Introduction
Population heterogeneity has been recognised as a fundamental determinant of epidemic dynamics. Heterogeneities may be due to age [3, 21] or spatial location [22], and in this context, the structure of contact matrices [17] or connections between nodes [10] play an important role; heterogeneity in the number of contacts has been considered especially in the context of sexually-transmitted infections [13], in which case the type of mixing (assortative or proportionate) significantly affects the epidemic dynamics. A different kind of heterogeneity, differences in susceptibility, i.e. in the probability of getting infected when contacted by an infectious individual, has been proposed by several authors [7] during the Covid epidemic and systematically developed by [18] and, more recently, by [11] (see also [6] and [14]). A somewhat similar model had been studied by [12], who analysed the long-term dynamics of an endemic SIR model in which the population is divided into a discrete number of groups differing in susceptibility to infection. Different approaches to incorporate heterogeneity have been proposed by [20] and by [5].
[11] developed the theory for a general type of epidemic in which infectiousness depends, in an arbitrary way, on the time since infection, and the model is set as a renewal equation; later, they specialise the resulting equation for the case in which the model can be described by a system of ODEs. Here, we consider only the case of SEIR models, which are widely used to describe the dynamics of influenza epidemics, as well as several other infectious diseases.
The paper is organised as follows. In Section 2, we derive the equations for this special case, following [11], under two alternative assumptions on the relationship between susceptibility and infectiousness. By concentrating on this specific setting, the derivation becomes more transparent, even though the resulting systems have already been presented in [11]. In Section 3 we study the epidemic final size for the resulting models; there we show, similarly to [18], that the presence of heterogeneity always decreases the epidemic final size relative to a homogeneous model with the same value of $\mathcal{R}_{0}$, under either assumption on the relation between susceptibility and infectiousness. In Section 4, we specialise the equations to the case where susceptibility heterogeneity follows either a gamma [11] or an extended beta distribution. In Section 5, we prove that, in the case of gamma distributions, a greater variance in susceptibility lowers the attack ratio, and that, for the same value of $\mathcal{R}_{0}$, the attack ratio is lower when infectiousness is proportional to susceptibility than when it is independent; we also compare numerically the final size obtained with the gamma distribution with that resulting from beta distributions with the same variance. Finally, in Section 6, we compare models with and without susceptibility heterogeneity by fitting them to Italian seasonal influenza data from 2014-15 to 2023-24.
@@ 2 Model construction
The feature $x$ could depend on genetic factors, immunological history, or possibly behaviour. Still, it is assumed that it is fixed, at least as long as an epidemic outbreak is concerned, and that its distribution in the population has density $\phi(x)$ (a more general distribution, allowing also for a discrete part, is considered in [11], but here we stick to the simplest case).
Following [11], we can set $a(x)=x$ (one needs only to redefine the feature) and assume that its mean is 1, i.e., $\int_{0}^{\infty}x\phi(x)\,\textnormal{d}x=1$, by redefining $F(t)$ appropriately.
@@
The subsequent elaborations depend on the choice of the function $c(x)$; following [11], we consider the cases $c(x)\equiv\beta$ (infectiousness independent of susceptibility) and $c(x)=\beta x$ (perfect correlation between susceptibility and infectiousness).
@@
While (8) is a closed system, it may be convenient to also consider the variables $E(t)$ and $V(t)$, in order to compare the model results with data. One has
@@
Although the function $\Phi_{k}$ cannot generally be explicitly computed (a special case is the gamma distribution for $x$, see Section 4.1), there is no problem in numerically integrating system (7).
@@ 3 Basic properties
In the case of correlated infectiousness and susceptibilities, $\Phi_{2}(1)$ involves the variance of the susceptibility $X$, since $\Phi_{2}(1)=1+{Var(x)}$. The resulting formula is similar to that obtained with the proportionate mixing in the case of varying contact rates [2].
@@
The proof can be found in Appendix A. In biological terms, Theorem 3.1 shows that with any distribution of heterogeneity in susceptibility, and with both models, the susceptible fraction at the end of an epidemic is larger than with homogeneous susceptibility; in other words, the attack ratio (or epidemic final size) will be smaller. This is similar to the results obtained by [18].
@@ 4.1 The case of gamma-distributed susceptibility
Again following [11], we assume that $X$ follows the gamma distribution with mean 1 and variance $1/p$, i.e.,
@@
Figure 1: $p=0.005,R_{0}=\beta\cdot\tfrac{1}{\gamma}=\tfrac{1}{7}=$
Figure 2: $p=0.6,R_{0}=$
Figure 3: $p=5,R_{0}=$
Can the peak of the green curve be this overwhelmingly high? The flow from $S$ to $E$ is a bit confusing.
@@ 4.2 Beta-distributed susceptibility
Here we assume that $X$ follows the Beta distribution with shape parameters $a$ and $b$, scaled by a factor $L>1$, so that its support becomes $(0,L)$ (we will sometimes refer to this as an extended Beta distribution). Then
@@
Notice that, for a given value of $a$, if one lets $V$ increase to $1/a$ (the largest possible value) while using (19), the distribution converges to a Gamma(a)(a).
@@
With this parametrization, if $L$ is fixed, and $V$ increases to L1L-1 (the largest possible value) while using (20), the distribution converges to a bimodal distribution where $X=L$ with probability $1/L$ and $X=0$ with probability (L1)/L(L-1)/L.
@@
This expression can be rewritten in terms of Kummer’s function $M(a,b,z)$ [1, p. 505], also known as the hypergeometric confluent function $_{1}F_{1}(a,b,z)$ whose numerical computation is provided in several software packages (e.g. Wolfram-Alpha, Python, or Matlab), making computations much faster. Precisely, one has
@@ 4.3 Numerical illustrations
Figure 5: Plots showing proportion of $E(t)+I(t)$ (above) and $R(t)$ (below) for differently distributed susceptibility over time (all with $\mathcal{R}_{0}=2$). On the left panels, plots are shown with infectiousness independent of susceptibility; on the right panels, one can see plots with directly correlated infectiousness and susceptibility. The black curve represents the homogeneous SEIR model, coloured solid curves represent gamma distributions with different variance (see legend), coloured dashed curves represent Beta-distributed heterogeneity, with the same variance $(Var(X)=1/3)$ as the light-blue, solid curve. Note that $U(0,2)$ is the uniform distribution in $(0,2)$, which is the same as the extended Beta with $a=b=1$.
It can be seen that the homogeneous SEIR model yields the largest infection peak and the largest final size. Moreover, as the variance of the gamma distribution increases, the final size systematically decreases, both when susceptibility and infectiousness are independent and when they are correlated. This dependence of the final size on the variance will be proved in the next section. One can also observe that different Beta distributions with the same variance produce very similar final epidemic sizes, both among themselves and relative to the gamma distribution with the same variance. In this sense, the epidemic dynamics appear to depend primarily on the variance of the susceptibility distribution rather than on its specific shape: although the three distributions shown in the right panel of Figure 4 differ markedly in form, they generate very similar trajectories and final epidemic sizes.
5 Final size equation for the Gamma and Beta distributions
@@ 5.1 The case of independent infectiousness and susceptibility
Since the variance of the gamma distribution decreases with increasing $p$, this Theorem shows that the epidemic final size $(1-S_{\infty}(p)$ decreases with increasing variance. This result had been essentially proved by [18] with completely different notation and methods.
@@ 5.2 The case of correlated susceptibility and infectiousness
is constant along the solutions to (15). Moreover, it is not difficult to show that necessarily
@@
$R_{0}=1.5$ $R_{0}=3$
Refer to caption Refer to caption
Figure 6: Left panel: Attack ratio at the end of an epidemic for systems (7) and (8)–(LABEL:sys_bar2b), plotted as a function of the variance with $R_{0}=1.5$; curves are shown when susceptibility $X$ follows the gamma distribution, or the Beta distributions with either the parameter choices in (19) or (20) and when $X$ is either equal to 0 or to $V+1$ (bimodal distribution). Each distribution is shown (with the same line type) both for independent (the higher curve) and for correlated (the lower curve) infectiousness. Right panel: Same with $R_{0}=3$; only the curves corresponding to the Gamma (both with independent anc correlated infectiousness) and bimodal distributions are shown.
@@
We considered the final size also when susceptibility follows the (extended) Beta distribution, but this was possible only numerically, as shown below. We start by considering the limiting case, introduced in the previous Section, of a bimodal distribution, in which susceptibility is either equal to 0 (with probability (L1)/L(L-1)/L) or to $L$ (with probability $1/L$). Note that the probabilities have been set to have average susceptibility equal to $1$ and that their variance is equal to L1L-1.
This case actually corresponds to the standard epidemic model, where only a fraction $1/L$ is initially susceptible (while the rest of the population is immune) and one would set $S_{0}=1/L$, and the contact parameter would be defined as $\tilde{\beta}=\beta L$, and the basic reproduction number $\tilde{\mathcal{R}}_{0}=\tilde{\beta}/\gamma$, while what here is defined as $\mathcal{R}_{0}=\beta/\gamma$ would be the effective reproduction number. One can then use the standard formula for $\bar{x}=S_{\infty}/S_{0}$ that is obtained (if $\mathcal{R}_{0}=\tilde{\mathcal{R}}_{0}S_{0}>1$ as the only solution in $(0,1)$ of
$F(x):=\log(x)+\mathcal{R}_{0}(1-x)=0.$ (25)
Hence, the attack ratio is $S_{0}(1-\bar{x})=\dfrac{1-\bar{x}}{V+1}$.
Note, finally, that in this case one cannot distinguish between independent and correlated infectiousness, since only those who are susceptible can be infected.
Comparing this case with the case where the susceptibility follows a Gamma-distribution, we have the following
Proposition 5.4.
$\dfrac{1-\bar{x}}{V+1}<1-S_{\infty}(p)$ where $\bar{x}$ is solution of (25), $S_{\infty}(p)$ solution of (23), and $p=1/V$.
In words, the attack ratio found when susceptibility is bimodal is lower than that obtained when susceptibility is gamma-distributed with the same variance and the same value of $\mathcal{R}_{0}$.
We complete the analysis with numerical computations of the attack ratio. In Figure 6, we show how the final attack ratio (the fraction of the population infected during an epidemic outbreak) depends on the variance for some distributions of susceptibility. In the left panel ($\mathcal{R}_{0}=1.5$), we show the results for the gamma distribution and for the Beta distribution under the parametrisations (19) and (20) with either independent or correlated infectiousness, and also for the bimodal distribution. The panel, beyond numerically confirming the results of Theorems 5.1 and 5.2, of Propositions 5.3 and 5.4, shows that the attack ratios obtained with the Beta distributions are intermediate between those with the gamma distribution and the bimodal case for the same value of the variance and the same type of infectiousness (independent or correlated to susceptibility). One can also see that, for the same distribution of susceptibility, the attack ratio when infectiousness is correlated to susceptibility is lower than when they are independent, extending the result proved for the gamma distribution in Proposition 5.3. Furthermore, in the limiting cases where the Beta distribution tends either to the Gamma or to the bimodal case, the attack ratio converges, as expected, to the corresponding values. Finally, one sees that, for all distributions, the attack ratios obtained when infectiousness is correlated to susceptibility are lower than for the bimodal case.
In the right panel, we show how the final attack ratio depends on the variance of susceptibility when $\mathcal{R}_{0}=3$, only for the gamma distribution (both with independent and correlated infectiousness) and for the bimodal case. It can be seen that the final observation made when $\mathcal{R}_{0}=1.5$ does not hold: in this case, the attack ratio with bimodal susceptibility is lower than with correlated infectiousness. All other observations concerning Beta distributions (the attack ratios being intermediate between the corresponding Gamma and bimodal, converging to the expected limiting case, being lower with correlated than with independent infectiousness) are instead confirmed. This can be seen in the Figures shown in the Supplementary Material (Section B), since we avoided showing the corresponding lines in the right panel, which would have become very messy with all the lines.
Overall, one can say that the final epidemic size obtained from the Beta distribution is close to that produced by a gamma distribution with the same variance and the same choice of either independent or correlated infectiousness, though some differences can be seen. When comparing Beta and gamma distributions, or Beta distributions with different parametrisations, it is possible to find pairs of distributions for which the one with the lower variance also has a lower attack ratio, in line with the observation of [18] that mean and variance do not fully characterise the behaviour of heterogeneous models. Nevertheless, within each fixed Beta parametrisation, increasing the variance consistently leads to a lower attack ratio.
@@ 6.1 Data description
The data used in this study were obtained from RespiVirNet [15], which provides weekly, age-stratified influenza-like illness (ILI) incidence rates from the national sentinel surveillance system. Raw records include ILI aggregated case counts (and number of patients in the sentinel cohorts) from the 2014-15 season onwards, reported by age groups (0–4, 5–14, 15–64, 65 and above) and week (from Week 46 to Week 17 of the subsequent year). Furthermore, data include, for each week, the fraction of virological samples from the sentinel network that are positive to the different influenza strains; these are age-stratified only in recent years, thus, for consistency, we did not use this information.
@@
  • $ur_{a}$ is the estimate of age-dependent underreporting ratio [23].
The influenza seasons considered range from 2014-15 to 2023-24. However, the seasons 2020-21, 2021-22, and 2022-23 have been excluded from the analysis since the lockdowns and other protective measures strongly affected the dynamics (no influenza cases were recorded in 2020-21, very few in 2021-22, while 2022-23 had an anomalous pattern, incompatible with a SEIR model with constant parameters).
@@ 6.2 Fitting methods
For the model (14) with gamma-distributed susceptibility, we considered as unknown parameters $\beta$, $p$, and the initial condition $I_{0}$ assuming for simplicity that $E(0)=I(0)=I_{0}$, where 0 is the first week of each epidemic season; at $t=0$ everybody else was assumed to be susceptible, as nobody was considered completely immune to the circulating strain at the season start. In the case of the homogeneous SEIR model (formally, it can be seen as model (14) with $p=\infty$), we instead fitted $\beta$ and the initial conditions $E(0)=I(0)$ and $R(0)$. For both models, we set $\alpha=\gamma=4$ (weeks-1), corresponding to an average length of generation time of 3.5 days [25].